OpenStudy (anonymous):

Can anyone help me out with this? Find the inverse of the function. f(x) = 6x3 + 3

OpenStudy (aravindg):

Put y=6x3+3 Then solve for x

OpenStudy (anonymous):

How would i solve for x?

OpenStudy (anonymous):

jst find the value of x in the terms of y where y = f(x) thus \[y = 6x ^{3} + 3\] and thus u will get \[x= \sqrt[3]{\frac{ y }{ 6 } - \frac{ 1 }{ 2 }}\]

OpenStudy (jhannybean):

Then don't you have to write back in terms of f^(-1) x such as \[f ^{-1}(x)=\sqrt[3]{\frac{ x }{ 6 }-\frac{1}{2}}\]